# A note for double Hölder regularity of the hydrodynamic pressure for weak solutions of Euler equations

### v3

@@ Abstract.
We give an elementary proof for the interior double Hölder regularity of the hydrodynamic pressure for weak solutions of the Euler Equations in a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$; $d\geq 3$. That is, for velocity $u\in C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ with some $20<\gamma<1/$ 2, we show that the pressure $p\in C^{0,2\gamma}_{\rm int}(\Omega)$. This is motivated by the studies of turbulence and anomalous dissipation in mathematical hydrodynamics and, recently, has been established in [L. De Rosa, M. Latocca, and G. Stefani, Int. Math. Res. Not. 2024.3 (2024), 2511–2560] over $C^{2,1}$-domains by means of pseudodifferential calculus. Our approach involves only standard elliptic PDE techniques, and relies on a variant of the modified pressure introduced in [C. W. Bardos, D. W. Boutros, and E. S. Titi, Hölder regularity of the pressure for weak solutions of the 3D Euler equations in bounded domains, Arch. Rational Mech. Anal. 249 (2025), 28] and the potential estimates in [L. Silvestre, unpublished notes]. The key novel ingredient of our proof is the introduction of two cutoff functions whose localisation parameters are carefully chosen as a power of the distance to ${\partial\Omega}$.
@@ 1.1. Main Theorem
Let $(u,p)$ be a weak solution of the Euler equation (1.1) in the bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$ for $d\geq 3$. Assume that $u$ is of Hölder regularity $C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ in the spatial variable for some $20<\gamma<1/$ 2. Then the hydrodynamic pressure $p$ is locally of double Hölder regularity $C^{0,2\gamma}_{\rm loc}(\Omega)$ in the spatial variable. Moreover, for any subdomain ${\widetilde{\Omega}}\Subset\Omega$ and any $x_{1},x_{2}\in{\widetilde{\Omega}}$,
@@
with the constant $C=C\left(d,\gamma,{\widetilde{\Omega}}\right)\cdot\|u\|_{C^{0,\gamma}(\Omega)}^{2}$, where $C\left(d,\gamma,{\widetilde{\Omega}}\right)$ depends only on $d$, $\gamma$, the $C^{2}$-geometry of $\Omega$, and the distance from ${\widetilde{\Omega}}$ to $\partial\Omega$.
@@
In the above theorem, more precisely, one may replace the “$C^{2}$-geometry of $\Omega$” with the following geometrical data: the $C^{0}$-norm of the second fundamental form of ${\partial\Omega}$, the intrinsic diameter of $\Omega$ (i.e., the infimum of the length of $C^{1}$-curves connecting any two points inside $\Omega$), and the injectivity radius of $\Omega$. If a constant $C_{\rm geom}$ depends on these data only, we shall say that $C_{\rm geom}$ is a geometrical constant.
@@
Bardos and Titi initiated in [3] the project of extending the Onsager conjecture and related results to bounded domains $\Omega\subset\mathbb{R}^{d}$. The Hölder regularity of the hydrodynamic pressure $p$ in the Euler equations (1.1) plays a fundamental role in the study of anomalous dissipation on bounded domains [3, 5, 19, 20]. Assuming $u\in C^{0,\gamma}(\Omega)$ with $10<\gamma<$ 1 and $\Omega$ is a bounded $C^{3}$-domain, Bardos and Titi obtained that $p\in C^{0,\gamma}_{\rm int}(\Omega)$ for $\Omega\subset\mathbb{R}^{2}$ in [4] and, together with Boutros, proved the same result for $\Omega\subset\mathbb{R}^{3}$ in [2]. Then, utilising among others the techniques developed in [7, 8], De Rosa, Latocca and Stefani [13] established the almost double Hölder regularity in the sense that $p\in C^{1,\min(\alpha,2\gamma-1)}(\Omega)$ for $11/2<\gamma<$ 1 on $C^{2,\alpha}$-domains, and $p\in C^{2\gamma-\varepsilon}(\Omega)$ with arbitrarily small $\varepsilon>0$ for $20<\gamma<1/$ 2 on $C^{3,\alpha}$-domains. The same authors later in [14] extended the double Hölder regularity $p\in C^{2\gamma}(\Omega)$ to the index range $20<\gamma<1/$ 2 on $C^{2,1}$-domains. In addition, for the critical index $\gamma=1/2$, it is proved in [14, Proposition 2.5] that $p\in C^{1}_{\star}(\Omega)$, the Calderón–Zygmund space. This improves the log-Lipschitz regularity of $p$ obtained earlier in Constantin [9]. The interior double Hölder regularity of $p$ on $C^{4}$-domains in $\mathbb{R}^{3}$ has also been proved in [2].
The approach in the recent work [14] by De Rosa, Latocca and Stefani involves a delicate application of pseudodifferential calculus and the Littlewood–Paley theory. Here, on a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d\geq 3}$, we prove the interior double Hölder regularity $p\in C^{2\gamma}_{\rm loc}(\Omega)$ for $20<\gamma<1/$ 2, utilising only elementary techniques of (Neumann) Green functions and integration by parts. Moreover, we give explicit bounds on the blowup rate for $\frac{|p(x)-p(y)|}{|x-y|^{2\gamma}}$ as $x,y$ approach the boundary ${\partial\Omega}$. The issue of double Hölder regularity of $p$ up to the boundary might have remained to be clarified, as indicated in [2, p.8, at the end of § 1.3].
Our arguments make essential use of ideas from Silvestre [22] and Bardos–Boutros–Titi [2]. As in [2] (with an extra correction term), we introduce a modified pressure Weierstrass-$\wp$ (see §2 for details), which equals
@@
on ${\partial\Omega}$ and equals $p$ outside the $(2\delta)$-collar of the boundary, with $\delta>0$ being a small parameter to be specified. Here, ${\rm II}$ is the second fundamental form of ${\partial\Omega}$, $u^{\top}$ is the horizontal components of $u$, and $y_{\star}$ is the unique point on ${\partial\Omega}$ such that $|y-y_{\star}|={\rm dist}(y,{\partial\Omega})$. The construction of Weierstrass-$\wp$ makes use of a cutoff function ${\varphi^{\delta}}$ localised in the $(2\delta)$-collar of ${\partial\Omega}$.
Rather than working directly with Weierstrass-$\wp$, we introduce a second cutoff function ${\eta^{\delta}}$ (see (3.4)) which equals zero in the $\delta/2$-collar of ${\partial\Omega}$ and equals one outside the $\delta$-collar. Thus, by working with the elliptic PDE for ${\eta^{\delta}}\wp$ — which is localised to the interior of $\Omega$ — we circumvent the issues caused by highly irregular boundary values (see Remark 2.1 and [2, Section 8]). The cost to pay is that the singular terms $\left|\nabla{\eta^{\delta}}\right|\approx\delta^{-1}$ and $\left|D^{2}{\eta^{\delta}}\right|\approx\delta^{-2}$ enter our estimates. We overcome this issue by carefully adjusting the “boundary layers” induced by cutoff functions. For reasons that will become transparent along the proof, when estimating $|\wp(x_{1})-\wp(x_{2})|$ for $x_{1}$, $x_{2}$ sufficiently close to each other and $|x_{1}-x_{2}|\approx{\rm dist}\left(\frac{x_{1}+x_{2}}{2},{\partial\Omega}\right)$, we choose
@@ 1.2. Notations
A continuous function $u:{\widetilde{\Omega}}\to\mathbb{R}^{\ell}$ is $\alpha$-Hölder ($10<\alpha<$ 1) if the following seminorm is finite:
@@
A function $u:\Omega\to\mathbb{R}^{\ell}$ is in $C^{0,\alpha}_{\rm loc}(\Omega)$ if its restriction on ${\widetilde{\Omega}}$ is in $C^{0,\alpha}\left({\widetilde{\Omega}}\right)$ for any ${\widetilde{\Omega}}\Subset\Omega$.
@@
Starting from any $x\in{\partial\Omega}$, one may flow the point $x$ by the inward unit normal $-\nu(x)$ for some time $\tau(x)$. The supremum of those numbers $\tau>0$ such that if $\tau(x)\leq\tau$ for all $x\in{\partial\Omega}$ then the image of ${\partial\Omega}$ under the flow has no self-intersections is the injectivity radius of $\Omega$, denoted as ${\rm injrad}(\Omega)$. For the bounded $C^{2}$-domain $\Omega$, one has ${\rm injrad}(\Omega)>0$. The distance function $y\mapsto{\rm dist}(y,{\partial\Omega})$ lies in $C^{2}\left(\Omega\setminus\overline{\Omega_{\delta}}\right)$ for any $\delta\in\left]0,{\rm injrad}(\Omega)\right[$.
@@ 1.3. Organisation
In §2 we define the modified pressure Weierstrass-$\wp$, which involves the first cutoff function ${\varphi^{\delta}}$ and is a variant of the one introduced in Bardos, Boutros, and Titi [2].
The main result, Theorem 1.1, will be proved in §3. The novel tool for our proof, namely that the second cutoff function ${\eta^{\delta}}$, is elaborated in §3.1. Estimates for Weierstrass-$\wp$ occupy §§3.2 &\&& 3.3. Finally, the proof of Theorem 1.1 is concluded in §3.4.
@@ 2. The modified pressure
In this section, motivated by Bardos, Boutros and Titi [2, (2.17)], we introduce the modified pressure Weierstrass-$\wp$ from the problem (1.1).
@@
$\partial_{\nu}p=\nabla{\widetilde{\nu}}:(u\otimes u)-\partial_{\nu}\left(u\cdot{\widetilde{\nu}}\right)^{2}-\partial_{\tau}[(u\cdot\tau)(u\cdot{\widetilde{\nu}})]-\partial_{t}(u\cdot{\widetilde{\nu}}),$
@@
To proceed, let us further introduce a second modified pressure Weierstrass-$\wp$, which satisfies the homogeneous Neumann boundary condition and hence admits a representation formula via the Neumann Green function (see (3.2) and the ensuing lines). For this purpose, we set
@@
$\partial_{{\widetilde{\nu}}}{\mathcal{R}}(y)$ =$=\left\{\partial_{{\widetilde{\nu}}}{\rm dist}(y,{\partial\Omega}){\varphi^{\delta}}(y)+{\rm dist}(y,{\partial\Omega})\partial_{{\widetilde{\nu}}}{\varphi^{\delta}}(y)\right\}\left[{\rm II}\left(u^{\top},u^{\top}\right)(y_{\star})\right]$
@@
for ${\mathcal{Q}}$ and ${\mathcal{R}}$ in (2.4) and (2.6), respectively, we obtain the homogeneous Neumann boundary problem for Weierstrass-$\wp$:
@@
Equation (2.8) serves as the starting point of all our subsequent developments. It is straightforward to check that (2.5) and (2.8) simultaneously satisfy the compatibility condition for the Neumann problem of the Poisson equation. For simplicity, we refer to Weierstrass-$\wp$ as the modified pressure from now on. Without loss of generality, we assume the normalisation condition:
@@
Indeed, the solutions Weierstrass-$\wp$ may differ from the one with average zero by addition of some function whose gradient is a harmonic vector field, which is nontrivial on non-simply-connected $\Omega$. But such a function is in $C^{\infty}(\Omega)$ when $\Omega$ is a $C^{2}$-domain, hence does not affect the double Hölder regularity of Weierstrass-$\wp$.
@@
In this section, we prove our main Theorem 1.1 by establishing the bound for Weierstrass-$\wp$ in the $C^{0,2\gamma}(\Omega)$-norm from the homogeneous Neumann problem (2.8). The key new idea is to introduce, other than ${\varphi^{\delta}}$ in (2.7), a second cutoff function localised in $\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}$.
@@ 3.1. The second cutoff function
$\begin{cases}\text{the segment }[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}\text{ and }c_{\Omega}\delta\leq|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta},\\ \text{where $\beta:=\frac{d-2}{d-2+2\gamma}\in]0,1[$ and $c_{\Omega}>0$ is a geometrical constant.}\end{cases}$ (3.1)
@@ 3.2. Estimate for $I_{1}$
  • Weierstrass-$\wp$ satisfies the Neumann problem (2.8) of the form $\Delta\wp=D^{2}F$, where $F\in C^{0,\gamma}(\Omega)$ is quadratic in $u$. Thus, away from the boundary ${\partial\Omega}$, we have $\wp\in C^{0,\gamma}(\Omega_{\delta/2})$.
@@ 2. Denote
:=$:=\left\{\int_{\Omega_{\rm near}}+\int_{\Omega_{\rm far}}\right\}\big{(}u^{i}(y)-u^{i}(x_{1})\big{)}\big{(}u^{j}(y)-u^{j}(x_{2})\big{)}\partial_{i}\partial_{j}\left({\eta^{\delta}}{\psi_{x_{1},x_{2}}}\right)(y)\,{\rm d}y.$
@@
For any $y\in\Omega_{\rm near}$, we have $|y-\bar{x}|\leq 5\rho\leq 5c_{\Omega}\delta^{\beta}$. (Recall from condition (3.1) that $\rho=|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$ for $\beta=\frac{d-2}{d-2+2\gamma}$, and that the segment $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$.) The triangle inequality yields that so $y\in\overline{\Omega_{5\delta^{\beta}}}$. In particular, $y$ is not in the annulus $\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}$, which contains the support of derivatives of ${\eta^{\delta}}$ (see (3.4)). We thus have the identity
@@
×$\qquad\qquad\qquad\qquad\times\Big{\{}\left|\partial_{i}\partial_{j}{\mathcal{K}_{N}}(y-x_{1})\right|+\left|\partial_{i}\partial_{j}{\mathcal{K}_{N}}(y-x_{2})\right|\Big{\}}\,{\rm d}y$
@@
×$\qquad\times\left\{\frac{\rho}{|y-\bar{x}|^{d+1}}+\left(\frac{\rho}{\delta|y-\bar{x}|^{d}}+\frac{\rho}{\delta^{2}|y-\bar{x}|^{d-1}}\right){\mathbf{1}}_{{\Omega_{\delta/2}\setminus\Omega_{\delta}}}\right\}\,{\rm d}y.$ (3.13)
@@
The integral $\int_{\Omega_{\rm far}}|y-x_{1}|^{\gamma}|y-x_{2}|^{\gamma}\frac{\rho}{|y-\bar{x}|^{d+1}}\,{\rm d}y$ can be treated similarly as in [22, Section 1.2]. In view of the triangle inequality and $10<\gamma<$ 1, we have
@@
From (3.1) and Lemmas 3.1 &\&& 3.2, we deduce
@@
provided that the condition (3.1) is verified. That is, the above estimate holds when $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$ and $c_{\Omega}\delta\leq|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$, where $10<\beta<$ 1 is a fixed number given in (3.8), and $c_{\Omega}>0$ is geometrical. The constant may be chosen as $C=C(d,\gamma)\cdot C_{\rm geom}\cdot\|u\|_{C^{0,\gamma}(\Omega)}^{2}.$ Then, in view of the definition of Weierstrass-$\wp$ (see (2.7)), we have =pWeierstrass-$\wp=p$ in $\Omega\setminus{\bf supp}({\varphi^{\delta}})$, where $p$ is the pressure we need to estimate. But ${\bf supp}({\varphi^{\delta}})\subset\Omega\setminus\Omega_{2\delta}$ while $[x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$, so $x_{1},x_{2}\notin{\bf supp}({\varphi^{\delta}})$. We thus obtain the desired estimate (1.2) in Theorem 1.1, for those $x_{1}$, $x_{2}$ satisfying the condition (3.1).
@@
leads to $\kappa>12\delta^{\beta}$. On the other hand, $\rho<\left(c_{\Omega}\right)^{-\frac{\beta}{1-\beta}}$ implies that $\rho\leq\delta^{\beta}$. Hence, we obtain $[x_{1},x_{2}]\subset\Omega_{10\delta^{\beta}}$, by replacing $c_{\Omega}$ with a larger geometric constant if necessary.

### v4

@@ Abstract.
We give an elementary proof for the interior double Hölder regularity of the hydrodynamic pressure for weak solutions of the Euler Equations in a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$; $d\geq 3$. That is, for velocity $u\in C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ with some $0<\gamma<1/2$, we show that the pressure $p\in C^{0,2\gamma}_{\rm int}(\Omega)$. This is motivated by the studies of turbulence and anomalous dissipation in mathematical hydrodynamics and, recently, has been established in [L. De Rosa, M. Latocca, and G. Stefani, Int. Math. Res. Not. 2024.3 (2024), 2511–2560] over $C^{2,1}$-domains by means of pseudodifferential calculus. Our approach involves only standard elliptic PDE techniques, and relies on a variant of the modified pressure introduced in [C. W. Bardos, D. W. Boutros, and E. S. Titi, Hölder regularity of the pressure for weak solutions of the 3D Euler equations in bounded domains, Arch. Rational Mech. Anal. 249 (2025), 28] and the potential estimates in [L. Silvestre, unpublished notes]. The key novel ingredient of our proof is the introduction of two cutoff functions whose localisation parameters are carefully chosen as a power of the distance to $\partial\Omega$}.
@@ 1.1. Main Theorem
Let $(u,p)$ be a weak solution of the Euler equation (1.1) in the bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$ for $d\geq 3$. Assume that $u$ is of Hölder regularity $C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ in the spatial variable for some $0<\gamma<1/2$. Then the hydrodynamic pressure $p$ is locally of double Hölder regularity $C^{0,2\gamma}_{\rm loc}(\Omega)$ in the spatial variable. Moreover, for any subdomain ${\widetilde{\Omega}}\Subset\Omega$ and any $x_{1},x_{2}\in{\widetilde{\Omega}}$,
@@
[…]
@@
In the above theorem, more precisely, one may replace the $C^{2}$-geometry of $\Omega$” with the following geometrical data: the $C^{0}$-norm of the second fundamental form of $\partial\Omega$}, the intrinsic diameter of $\Omega$ (i.e., the infimum of the length of $C^{1}$-curves connecting any two points inside $\Omega$), and the injectivity radius of $\Omega$. If a constant $C_{\rm geom}$ depends on these data only, we shall say that $C_{\rm geom}$ is a geometrical constant.
@@
Bardos and Titi initiated in [3] the project of extending the Onsager conjecture and related results to bounded domains $\Omega\subset\mathbb{R}^{d}$. The Hölder regularity of the hydrodynamic pressure $p$ in the Euler equations (1.1) plays a fundamental role in the study of anomalous dissipation on bounded domains [3, 5, 19, 20]. Assuming $u\in C^{0,\gamma}(\Omega)$ with $0<\gamma<1$ and $\Omega$ is a bounded $C^{3}$-domain, Bardos and Titi obtained that $p\in C^{0,\gamma}_{\rm int}(\Omega)$ for $\Omega\subset\mathbb{R}^{2}$ in [4] and, together with Boutros, proved the same result for $\Omega\subset\mathbb{R}^{3}$ in [2]. Then, utilising among others the techniques developed in [7, 8], De Rosa, Latocca and Stefani [13] established the almost double Hölder regularity in the sense that $p\in C^{1,\min(\alpha,2\gamma-1)}(\Omega)$ for $1/2<\gamma<1$ on $C^{2,\alpha}$-domains, and $p\in C^{2\gamma-\varepsilon}(\Omega)$ with arbitrarily small $\varepsilon>0$ for $0<\gamma<1/2$ on $C^{3,\alpha}$-domains. The same authors later in [14] extended the double Hölder regularity $p\in C^{2\gamma}(\Omega)$ to the index range $0<\gamma<1/2$ on $C^{2,1}$-domains. In addition, for the critical index $\gamma=1/2$, it is proved in [14, Proposition 2.5] that $p\in C^{1}_{\star}(\Omega)$, the Calderón–Zygmund space. This improves the log-Lipschitz regularity of $p$ obtained earlier in Constantin [9]. The interior double Hölder regularity of $p$ on $C^{4}$-domains in $\mathbb{R}^{3}$ has also been proved in [2].
The approach in the recent work [14] by De Rosa, Latocca and Stefani involves a delicate application of pseudodifferential calculus and the Littlewood–Paley theory. Here, on a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d\geq 3}$, we prove the interior double Hölder regularity $p\in C^{2\gamma}_{\rm loc}(\Omega)$ for $0<\gamma<1/2$, utilising only elementary techniques of (Neumann) Green functions and integration by parts. Moreover, we give explicit bounds on the blowup rate for $\frac{|p(x)-p(y)|}{|x-y|^{2\gamma}}$ as $x,y$ approach the boundary $\partial\Omega$}. The issue of double Hölder regularity of $p$ up to the boundary might have remained to be clarified, as indicated in [2, p.8, at the end of § 1.3].
Our arguments make essential use of ideas from Silvestre [22] and Bardos–Boutros–Titi [2]. As in [2] (with an extra correction term), we introduce a modified pressure $\wp$ (see §2 for details), which equals
@@
on $\partial\Omega$} and equals $p$ outside the $(2\delta)$-collar of the boundary, with $\delta>0$ being a small parameter to be specified. Here, $\rm II$} is the second fundamental form of $\partial\Omega$}, $u^{\top}$ is the horizontal components of $u$, and $y_{\star}$ is the unique point on $\partial\Omega$} such that $|y-y_{\star}|={\rm dist}(y,{\partial\Omega})$. The construction of $\wp$ makes use of a cutoff function $\varphi^{\delta}$} localised in the $(2\delta)$-collar of $\partial\Omega$}.
Rather than working directly with $\wp$, we introduce a second cutoff function $\eta^{\delta}$} (see (3.4)) which equals zero in the $\delta/2$-collar of $\partial\Omega$} and equals one outside the $\delta$-collar. Thus, by working with the elliptic PDE for $\eta^{\delta}$$\wp$ — which is localised to the interior of $\Omega$ — we circumvent the issues caused by highly irregular boundary values (see Remark 2.1 and [2, Section 8]). The cost to pay is that the singular terms $\left|\nabla{\eta^{\delta}}\right|\approx\delta^{-1}$ and $\left|D^{2}{\eta^{\delta}}\right|\approx\delta^{-2}$ enter our estimates. We overcome this issue by carefully adjusting the “boundary layers” induced by cutoff functions. For reasons that will become transparent along the proof, when estimating $|\wp(x_{1})-\wp(x_{2})|$ for $x_{1}$, $x_{2}$ sufficiently close to each other and $|x_{1}-x_{2}|\approx{\rm dist}\left(\frac{x_{1}+x_{2}}{2},{\partial\Omega}\right)$, we choose
@@ 1.2. Notations
A continuous function $u:{\widetilde{\Omega}}\to\mathbb{R}^{\ell}$ is $\alpha$-Hölder ($0<\alpha<1$) if the following seminorm is finite:
@@
[…]
@@
Starting from any $x\in{\partial\Omega}$, one may flow the point $x$ by the inward unit normal $\nu(x)$ for some time $\tau(x)$. The supremum of those numbers $\tau>0$ such that if $\tau(x)\leq\tau$ for all $x\in{\partial\Omega}$ then the image of $\partial\Omega$} under the flow has no self-intersections is the injectivity radius of $\Omega$, denoted as ${\rm injrad}(\Omega)$. For the bounded $C^{2}$-domain $\Omega$, one has ${\rm injrad}(\Omega)>0$. The distance function $y\mapsto{\rm dist}(y,{\partial\Omega})$ lies in $C^{2}\left(\Omega\setminus\overline{\Omega_{\delta}}\right)$ for any $\delta\in\left]0,{\rm injrad}(\Omega)\right[$.
@@ 1.3. Organisation
In §2 we define the modified pressure $\wp$, which involves the first cutoff function $\varphi^{\delta}$} and is a variant of the one introduced in Bardos, Boutros, and Titi [2].
The main result, Theorem 1.1, will be proved in §3. The novel tool for our proof, namely that the second cutoff function $\eta^{\delta}$}, is elaborated in §3.1. Estimates for $\wp$ occupy §§3.2 &\& 3.3. Finally, the proof of Theorem 1.1 is concluded in §3.4.
@@ 2. The modified pressure
In this section, motivated by Bardos, Boutros and Titi [2, (2.17)], we introduce the modified pressure $\wp$ from the problem (1.1).
@@
[…]
@@
To proceed, let us further introduce a second modified pressure $\wp$, which satisfies the homogeneous Neumann boundary condition and hence admits a representation formula via the Neumann Green function (see (3.2) and the ensuing lines). For this purpose, we set
@@
$\partial_{{\widetilde{\nu}}}{\mathcal{R}}(y)$ $=\left\{\partial_{{\widetilde{\nu}}}{\rm dist}(y,{\partial\Omega}){\varphi^{\delta}}(y)+{\rm dist}(y,{\partial\Omega})\partial_{{\widetilde{\nu}}}{\varphi^{\delta}}(y)\right\}\left[{\rm II}\left(u^{\top},u^{\top}\right)(y_{\star})\right]$
@@
for $\mathcal{Q}$} and $\mathcal{R}$} in (2.4) and (2.6), respectively, we obtain the homogeneous Neumann boundary problem for $\wp$:
@@
Equation (2.8) serves as the starting point of all our subsequent developments. It is straightforward to check that (2.5) and (2.8) simultaneously satisfy the compatibility condition for the Neumann problem of the Poisson equation. For simplicity, we refer to $\wp$ as the modified pressure from now on. Without loss of generality, we assume the normalisation condition:
@@
Indeed, the solutions $\wp$ may differ from the one with average zero by addition of some function whose gradient is a harmonic vector field, which is nontrivial on non-simply-connected $\Omega$. But such a function is in $C^{\infty}(\Omega)$ when $\Omega$ is a $C^{2}$-domain, hence does not affect the double Hölder regularity of $\wp$.
@@
In this section, we prove our main Theorem 1.1 by establishing the bound for $\wp$ in the $C^{0,2\gamma}(\Omega)$-norm from the homogeneous Neumann problem (2.8). The key new idea is to introduce, other than $\varphi^{\delta}$} in (2.7), a second cutoff function localised in $\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}$.
@@ 3.1. The second cutoff function
[…]
@@ 3.2. Estimate for $I_{1}$
  • $\wp$ satisfies the Neumann problem (2.8) of the form $\Delta\wp=D^{2}F$, where $F\in C^{0,\gamma}(\Omega)$ is quadratic in $u$. Thus, away from the boundary $\partial\Omega$}, we have $\wp\in C^{0,\gamma}(\Omega_{\delta/2})$.
@@ 2. Denote
$:=\left\{\int_{\Omega_{\rm near}}+\int_{\Omega_{\rm far}}\right\}\big(u^{i}(y)-u^{i}(x_{1})\big)\big(u^{j}(y)-u^{j}(x_{2})\big)\partial_{i}\partial_{j}\left({\eta^{\delta}}{\psi_{x_{1},x_{2}}}\right)(y)\,{\rm d}y.$
@@
For any $y\in\Omega_{\rm near}$, we have $|y-\bar{x}|\leq 5\rho\leq 5c_{\Omega}\delta^{\beta}$. (Recall from condition (3.1) that $\rho=|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$ for $\beta=\frac{d-2}{d-2+2\gamma}$, and that the segment $x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$.) The triangle inequality yields that so $y\in\overline{\Omega_{5\delta^{\beta}}}$. In particular, $y$ is not in the annulus $\overline{\Omega_{\delta/2}\setminus\Omega_{\delta}}$, which contains the support of derivatives of $\eta^{\delta}$} (see (3.4)). We thus have the identity
@@
$\qquad\qquad\qquad\qquad\times\Big\{\left|\partial_{i}\partial_{j}{\mathcal{K}_{N}}(y-x_{1})\right|+\left|\partial_{i}\partial_{j}{\mathcal{K}_{N}}(y-x_{2})\right|\Big\}\,{\rm d}y$
@@
$\qquad\times\left\{\frac{\rho}{|y-\bar{x}|^{d+1}}+\left(\frac{\rho}{\delta|y-\bar{x}|^{d}}+\frac{\rho}{\delta^{2}|y-\bar{x}|^{d-1}}\right){\mathbf{1}}_{{\Omega_{\delta/2}\setminus\Omega_{\delta}}}\right\}\,{\rm d}y.$ (3.13)
@@
The integral $\int_{\Omega_{\rm far}}|y-x_{1}|^{\gamma}|y-x_{2}|^{\gamma}\frac{\rho}{|y-\bar{x}|^{d+1}}\,{\rm d}y$ can be treated similarly as in [22, Section 1.2]. In view of the triangle inequality and $0<\gamma<1$, we have
@@
From (3.1) and Lemmas 3.1 &\& 3.2, we deduce
@@
provided that the condition (3.1) is verified. That is, the above estimate holds when $x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$ and $c_{\Omega}\delta\leq|x_{1}-x_{2}|\leq c_{\Omega}\delta^{\beta}$, where $0<\beta<1$ is a fixed number given in (3.8), and $c_{\Omega}>0$ is geometrical. The constant may be chosen as $C=C(d,\gamma)\cdot C_{\rm geom}\cdot\|u\|_{C^{0,\gamma}(\Omega)}^{2}.$ Then, in view of the definition of $\wp$ (see (2.7)), we have $\wp=p$ in $\Omega\setminus{\bf supp}({\varphi^{\delta}})$, where $p$ is the pressure we need to estimate. But ${\bf supp}({\varphi^{\delta}})\subset\Omega\setminus\Omega_{2\delta}$ while $x_{1},x_{2}]\subset\Omega_{10\cdot\delta^{\beta}}$, so $x_{1},x_{2}\notin{\bf supp}({\varphi^{\delta}})$. We thus obtain the desired estimate (1.2) in Theorem 1.1, for those $x_{1}$, $x_{2}$ satisfying the condition (3.1).
@@
leads to $\kappa>12\delta^{\beta}$. On the other hand, $\rho<\left(c_{\Omega}\right)^{-\frac{\beta}{1-\beta}}$ implies that $\rho\leq\delta^{\beta}$. Hence, we obtain $x_{1},x_{2}]\subset\Omega_{10\delta^{\beta}}$, by replacing $c_{\Omega}$ with a larger geometric constant if necessary.